# Open problems and perspectives on solving Friedrichs systems by Krylov approximation

### v1

@@ 4. Perspective on Krylov solvability of Friedrichs-type inverse problems
(4.2) $A^{*}Af\,=\,A^{*}g$
Here $A^{*}A$ is self-adjoint (see, e.g., [39, Theorem 12.11]), and $A^{*}g\in\mathrm{ran}\,A^{*}A\cap\mathcal{D}^{qa}(A^{*}A)$. These are precisely the conditions that allow to deduce, via Theorem 2.7, that there exists a unique solution
@@
On the other hand, one observes that the above solution $f$ also satisfies $Af=g$. Indeed, from $A^{*}Af=A^{*}g$ one sees that $Af-g\in\ker A^{*}$, whereas obviously $Af-g\in\mathrm{ran}\,A=(\ker A^{*})^{\perp}$ because $Af$ and $g$ belongs to $\mathrm{ran}\,A$. From the triviality of $\ker A^{*}\cap(\ker A^{*})^{\perp}$ one thus deduces that $Af=g$.

### v2

@@ 4. Perspective on Krylov solvability of Friedrichs-type inverse problems
(4.2) $A^{*}Af\,=\,A^{*}g\,.$
Here $A^{*}A$ is self-adjoint (see, e.g., [38, Theorem 12.11]), and $A^{*}g\in\mathrm{ran}\,A^{*}A\cap\mathcal{D}^{qa}(A^{*}A)$. These are precisely the conditions that allow to deduce, via Theorem 2.7, that there exists a unique solution
@@
On the other hand, one observes that the above solution $f$ also satisfies $Af=g$. Indeed, from $A^{*}Af=A^{*}g$ one sees that $Af-g\in\ker A^{*}$, whereas obviously $Af-g\in\mathrm{ran}\,A\subset(\ker A^{*})^{\perp}$ because $Af$ and $g$ belongs to $\mathrm{ran}\,A$. From the triviality of $\ker A^{*}\cap(\ker A^{*})^{\perp}$ one thus deduces that $Af=g$.
@@
5. Acknowledgements
This work is partially supported by the Italian National Institute for Higher Mathematics INdAM and the Alexander von Humboldt Foundation, Bonn.